A system of four(4) free particles each has a mass mi and initial position rand initial velocity v.

a) Describe your own four-particle system in detail.

b) Determine the position Rcm and velocity Vcm of the centre of mass of the system.

c) Determine the total linear momentum Ptot and the linear momentum Pcm of the centre of mass of the system.

d)Determine the total angular momentum Ltot and the linear momentum Pcm of the centre of mass of the system.



1
Expert's answer
2021-01-29T13:23:06-0500

As per the question,

It is given that the mass of the 4 particles are "m_i= m_1, m_2, m_3 , m_4"

Position of the particles are "r_i= r_1, r_2, r_3, r_4"

Velocity of the particles are "v_i=v_1, v_2, v_3, v_4"

i)


ii) "R_{cm}=\frac{m_1r_1+m_2r_2+m_3r_3+m_4r_4}{m_1+m_2+m_3+m_4}"


"v_{cm}=\frac{m_1v_1+m_2v_2+m_3v_3+m_4v_4}{m_1+m_2+m_3+m_4}"


iii) Total momentum of the system "P_{tot}=m_1v_1+m_2v_2+m_3v_3+m_4v_4"


Total momentum about the center of mass be "P_{com}=(m_1+m_2+m_3+m_4)v_{com}"


"=m_1v_1+m_2v_2+m_3v_3+m_4v_4"


iv) Angular momentum "L_{tot}=L_1+L_2+L_3+L_4"


"=m\overrightarrow{r_1}\times \overrightarrow{v_1}+m\overrightarrow{r_2}\times \overrightarrow{v_2}+m\overrightarrow{r_3}\times \overrightarrow{v_3}+m\overrightarrow{r_4}\times \overrightarrow{v_4}"


Angular momentum about the center of mass "\overrightarrow{L_{com}}=m_{total}\overrightarrow{r_{com}}\times \overrightarrow{v_{com}}"


"=(m_1+m_2+m_3+m_4)(\frac{m_1r_1+m_2r_2+m_3r_3+m_4r_4}{m_1+m_2+m_3+m_4})\times(\frac{m_1v_1+m_2v_2+m_3v_3+m_4v_4}{m_1+m_2+m_3+m_4})"

v) Total momentum about the center of mass and the total momentum always be same but angular momentum about the center of mass and the angular momentum about the origin will not be same.